3. La Besurta

TUCA DEL SALBAGUARDIA (SALBAGUARDIA PEAK) 2 hours and 30 min. 2,738 m. high

Bordering peak, 2,738 m. high, of easy but spectacular ascent. Follow the same route as in the Portillón de Benàs (consult the Portillón de Benàs excursion). This peak is on the West of the Portillón. From the same opening, departs a track to the left. Take it and go up zigzagging by the South slope of the peak to get to the top after 2 hours and a half (departing from the Pllan d’Están). The track is in very good condition but is quite ”aerial” (not suitable for anyone suffering from vertigo). The views are, again, gorgeous.